OpenStudy (anonymous):

Finding the theoretical yield For lab we mixed together copper sulfate and sodium carbonate. The formula equation is :CuSO4+Na2CO3--->CuCO3+Na2SO4. there is 0.5M for both Copper(II) sulfate and for sodium carbonate. Then we filtered it through filter paper. The next day we weighed the solids on the filter paper. The solids is the copper(II)carbonate, correct? I got 0.8 grams. Now I am stuck at how to find the percent yield. Would the actual yield be 0.8 grams? and what would be the theoretical yield? Which compound do I use? I also use the t chart method, how should it work?

OpenStudy (anonymous):

I also found the limiting reactant which is copper(II) sulfate

OpenStudy (anonymous):

hellooo?

OpenStudy (aaronq):

You're right, the actual yield is 0.8 grams. To find the theoretical yield you use the moles of the limiting reactant (because this limits how much you can make) to relate them to the product you're interested in. Set up a ratio of the moles and the stoichiometric coefficients: \(\sf \dfrac{moles~of~limiting~reactant}{limiting ~reactant's~coefficient}=\dfrac{moles~of ~product}{products~coefficient}\) plug in your values and solve. Next you need to convert to grams and the percent yield using the following equation: \(\sf Percent~yield=\dfrac{theoretical~yield-actual~yield}{theoretical~yield}*100\%\)

OpenStudy (aaronq):

Since you used molarity (0.5M), you need to know the volume of solution you used to find the moles of reactants.